OpenStudy (anonymous):

tell whether each equation is true,false, or open. 3(-6)+5=26-3

jimthompson5910 (jim_thompson5910):

3(-6)+5=26-3 -18+5=26-3 -13 = 23 So the equation is false

OpenStudy (anonymous):

thanks so much ur the only one thats helping me!!!

jimthompson5910 (jim_thompson5910):

np